OpenStudy (anonymous):

the function D(x) models the cumulative number of death fom a diease x years after 1984. Estimate the year when there were 95000 deaths D(x)=3680^2+5079x+5873 There approximately 95000 deaths in the year ?

OpenStudy (bahrom7893):

I think u meant: D(x)=3680x^2+5079x+5873

OpenStudy (bahrom7893):

Now to find the approximate year when there were 95000 deaths: D(x)=3680^2+5079x+5873=95000

OpenStudy (bahrom7893):

3680^2+5079x+5873=95000, solve for x

OpenStudy (bahrom7893):

3680^2+5079x-89127=0

OpenStudy (bahrom7893):

oh man I've been copying off your stuff so I'm also ending up with no x^2.. hold on

OpenStudy (bahrom7893):

Now to find the approximate year when there were 95000 deaths: D(x)=3680x^2+5079x+5873=95000 3680x^2+5079x+5873=95000, solve for x 3680x^2+5079x-89127=0 http://www.wolframalpha.com/input/?i=Solve%283680x%5E2%2B5079x-89127%3D0%29

OpenStudy (bahrom7893):

According to wolf, x1 = -5.7 and x2 = 4.3. Since we can't go back in time yet, our answer must be positive, therefore x = 4.3

OpenStudy (bahrom7893):

There approximately 95000 deaths in the year 1984+x There approximately 95000 deaths in the year 1984 + 4.3 There approximately 95000 deaths in the year 1988.3 or roughly 1988

OpenStudy (anonymous):

Thnak u for showing the steps, so that I can understand...

OpenStudy (bahrom7893):

Please click on Good Answer if I helped, I really want to get to the next level, thanks =)!

OpenStudy (anonymous):

I did and I always do for u, thanks....